### v1

# Moyun: A Diffusion-Based Model for Style-Specific Chinese Calligraphy Generation

Abstract
Although Chinese calligraphy generation has achieved style transfer, generating calligraphy by specifying the calligrapher, font, and character style remains challenging. To address this, we propose a new Chinese calligraphy generation model “Moyun” , which replaces the Unet in the Diffusion model with Vision Mamba and introduces the TripleLabel control mechanism to achieve controllable calligraphy generation. The model was tested on our large-scale dataset “Mobao” of over 1.9 million images, and the results demonstrate that “Moyun” can effectively control the generation process and produce calligraphy in the specified style. Even for calligraphy the calligrapher has not written, “Moyun” can generate calligraphy that matches the style of the calligrapher.
Index Terms:
Calligraphy, Diffusion, Mamba
I Introduction
Chinese calligraphy, with a history spanning over thousands of years, is a cherished cultural treasure of China which represents the artistic of Chinese characters handwriting. Chinese calligraphy is rich in variations. Chinese has tens of thousands of characters, each with a different meaning. Additionally, a single character can be written in various fonts, such as regular script, running script, cursive script, clerical script, seal script, and so on. Moreover, the writing of the same font differs between calligraphers, as shown in Figure 1a. The writing of the same font by the same calligrapher shows consistency, which we refer to as calligraphic style. The rich variations in Chinese calligraphy require an extended period of study for an average person to master. However, Chinese calligraphy is widely used, which has led people to explore the use of AI for generating Chinese calligraphy. Recently, GAN and Diffusion models are applied to Chinese calligraphy generation [1, 2, 3, 4], yielding impressive outcomes.
Figure 1: (a) shows the character “bai” (which means ”white” in Chinese) written in different fonts by various calligraphers. Each column represents a different font, and each row corresponds to a different calligrapher. (b) The first row shows calligraphy generated by Calliffusion, while the second row shows the ground truth. In the first column (san, regular script, Yan Zhenqing), the strokes in the ground truth are evenly spaced, but Calliffusion’s result is not. In the second column (ba, cursive script, Su Shi), Calliffusion’s output appears less stable than the ground truth from an aesthetic standpoint.
However, there are still some problems with the current generation models of Chinese calligraphy.
ZiGAN [2] is a seminal work that applies GAN to the field of Chinese calligraphy generation. ZiGAN constructed several small-scale datasets, each with consistent styles, and trained models separately on these datasets. As a result, the trained models gained the ability to transform standard printed characters into calligraphy [^1] In the following text, we use “Calligraphy” to refer to a single character image generated by the calligraphy generation model. with the style of the specific dataset used during training. However, it cannot generate calligraphy guided by specific calligraphers, fonts, and characters. Calliffusion [3] is the first Chinese calligraphy generation model based on diffusion model[5]. Calliffusion uses descriptive text involving “character, font, and calligrapher” to effectively guide the generation process. However, it relies on a BERT[6] model pre-trained on Chinese language dataset as the text encoder, which struggles to accurately understand domain-specific terms in calligraphy, such as the names of calligraphers. For example, a calligrapher’s name might be split into two tokens. Additionally, this approach introduces extra computational cost, which is unnecessary for calligraphy generation tasks. To better accomplish this task, we proposed a multilabel mechanism, where independent classification labels correspond to the calligrapher, font, and character, and these labels are combined to control the generation process.
Calliffusion uses an Diffusion architecture based on Unet[7], but Unet does not adequately fit the structural relationships between the strokes of the characters, which is shown in Figure 1b. Our model is based on the Vision Mamba[8], which processes images through patchify. Additionally, we incorporated the more efficient Mamba model[9]. Experiments demonstrate that our model provides a better fit to the structure between strokes.
The ZiGAN’s dataset[2] is relatively small, containing only 9 sub-datasets, with each sub-dataset consisting of 6,000 images of a single style. In contrast, we collected 1.9 million single-character calligraphy images with diverse styles and detailed annotations. Due to varying collection conditions and different noise distributions, traditional binarization methods performed poorly. To address this, we designed a new binarization method based on SAM[10], which achieved better results.
In summary, we have the following contributions:
  • We propose a Chinese calligraphy model called “Moyun”, which is capable of generating $256\times 256$ single character calligraphy images. Moyun produces stroke structures and brushstrokes that align with those of real calligraphy, achieving state-of-the-art (SOTA) quality in image generation.
  • We innovatively introduced a multilabel calligraphy generation control mechanism “TripleLabel”, which allows for generating characters with label of calligrapher, font and character. Additionally, it can generate characters that the calligrapher has never written before.
  • We constructed a large-scale, well-annotated Chinese calligraphy dataset “Mobao” containing more than 1.9 million high-quality binarized images using SAM.
II Related works
II-A Generative Adversarial Networks
Generative Adversarial Networks (GAN) [11] consists of a generator and a discriminator. The generator tries to produce data similar to real data, while the discriminator’s task is to distinguish between the generated data and real data. Research on GAN-based calligraphy generation has been ongoing for a long time, with the earliest work being zi2zi [4], an open-source project based on pix2pix [12]. After that, there were calliGAN [1], ZiGAN[2], and end-to-end[13] models. The most recent work is a style transfer model called CCST-GAN[14].
II-B Diffusion Model
The Diffusion model is a widely studied image generation model, with its structure initially established in DDPM [5] and DDIM [15], and the introduction of VAE[16] in LDM [17] further improving its application. DiT [18] brought the Transformer [19] architecture into the Diffusion model. Research on Diffusion-based calligraphy generation began relatively late. The earliest work is Calliffusion[3], which utilized a Unet-based Diffusion model. The most recent work is DP-Font[20], which employs a physical information neural network structure, achieving better control over character shapes.
III Method
Figure 2 shows the architecture of “Moyun”. “Moyun” is a diffusion model based on Vision Mamba[8], optimized with Mamba2[9]. Furthermore, we control the generation process by the TripleLabel mechanism.
Figure 2: “Moyun” architecture. The input latent noise is patched. The label is a combination of the calligrapher, font, and character. We used Mamba2-Replacement-Vision Mamba to process the patches.
III-A Preliminaries
Moyun is a diffusion model, so our training and inference follow the diffusion [5, 15] approach. The forward process gradually adds noise to the real data, specifically to our calligraphy image $x_{0}$. The diffusion process follows the equation as shown in (1) .
$\vspace{-5pt}q(x_{t}|x_{0})=\mathcal{N}\left(x_{t};\bar{\alpha}_{t}x_{0},(1-\bar{\alpha}_{t})\mathbf{I}\right)\vspace{-5pt}$ (1)
Furthermore, the noise at step $t$, $x_{t}$ can be obtained by equation
$x_{t}=\bar{\alpha}_{t}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon_{t}$ (2)
The diffusion model trains the reverse denoising process, and the denoising equation is as follows:
$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}(\mu_{\theta}(x_{t}),\Sigma_{\theta}(x_{t}))$ (3)
Our loss function is given by the following equation:
$\mathcal{L}_{\text{simple}}(\theta)=\|\epsilon_{\theta}(x_{t})-\epsilon_{t}\|^{2}\vspace{-5pt}$ (4)
III-B Block Design
To improve the model’s ability to capture the structure of calligraphy, we replaced the Unet in the Diffusion Model with Vision Mamba[8], thereby introducing the patchify mechanism. Additionally, we used the latest Mamba2[9] model instead of the first version of Mamba[21] in Vision Mamba. Thus, each patch corresponds to a small part of the calligraphic structure, and Mamba’s efficient contextual relational ability effectively matches the relationships between patches, thereby better learning the structure of the calligraphy.
The input image is first mapped into the latent space via a VAE[16], followed by patchify processing. For a square image with dimensions of $a\times a$, it is first mapped to a latent space of size $b\times b$ by the variational autoencoder (VAE). The latent representation is then divided into $n$ patches of size $p\times p$, where $n=\frac{b^{2}}{p^{2}}$, with $p$ being the side length of each patch and $b$ the side length of the latent space. In total, the entire Block is iterated $N$ times.
III-C TripleLabel control
To accommodate the triple-conditions control requirements of Chinese calligraphy generation, which include calligrapher, font, and character, we designed a TripleLabel control mechanism. Each calligrapher, font, and character is mapped to a unique class label, represented by a number, and each label is arbitrary. This allows the combination of labels to generate calligraphy that the calligrapher has not written before. In the model, the input label is transformed into the corresponding embedding vector through an embedding table. The three resulting embedding vectors are summed and used to control the generation process via a scale-shift mechanism[18].
This control mechanism significantly reduces computational cost compared to introducing a new text encoder. Furthermore, subsequent experiments demonstrated that this method of control is highly effective.
IV experiment
IV-A Dataset
Figure 3: (a) shows the directory structure of dataset “Mobao”, using the calligrapher “Fan Zhongyan” as an example. “Zu” and “Zi” only have single images, while “Bai” has multiple images. (b) demonstrates the binarization process, using the character “Zu” as an example to show the steps of selecting points, obtaining the mask, and resizing.
Dataset construct
To construct the large-scale annotated dataset “Mobao”, we scraped images from web and subject them to binarization processing. The collected calligraphy images include six fonts: regular script, running script, cursive script, clerical script, seal script, and seal carving. Each calligraphy image is annotated with the calligrapher, font, and character. We organized the data into a hierarchical folder structure of calligrapher-font-character, and each image within these categories was numbered. The complete processing workflow is illustrated in Figure 3.
To perform binarization, we employ SAM for image segmentation. Initially, Otsu binarization is applied to the image, and connected regions larger than 100 pixels are selected, which typically represent strokes in calligraphy. Within these connected regions, positive points are uniformly sampled using k-means. In addition, negative points are selected outside of these connected domains. The formula for selecting the number of points is as follows:
$\vspace{-5pt}cnt_{pos}=\max(1,\min(20,\frac{area_{regions}\times 100}{area_{total}})\vspace{-1pt}$ (5)
$\vspace{-1pt}cnt_{neg}=50$ (6)
These points are then provided to SAM as prompts with original calligraphy images, leading to a mask output by SAM, which we regard as the foreground. By filling the complement with white, we obtain the binarized image. Finally, a resizing process is conducted, in which the longer side of the image’s dimensions is stretched to 256 while maintaining the aspect ratio. Afterwards, it is centered and concatenated with a white background to yield a $256\times 256$ image.
dataset analyse
After processing the dataset, we conducted an analysis. We obtained a total of 1,929,393 images, including works from 6 fonts, 2,681 calligraphers and 4,660 characters. Some images without attributed calligraphers were categorized under the “anonymous” category.
To further evaluate our dataset, we conducted statistical analysis across the three dimensions of calligrapher, font, and character. The results confirm that our dataset is vast: the calligrapher with the most collected works, “Wang Xizhi”, has a total of 124,854 images; the font with the largest collection , running script, contains 668,040 images; and the character with the most instances, “shu” (book in chinese), appears in 11,949 images. This provides a rich source of data for the model to learn the structure and style of calligraphy.
However, our dataset has some issues. Nearly half of the calligraphers have fewer than ten collected works, and half of the characters have fewer than one hundred images. This indicates that our dataset exhibits a long-tail distribution.
IV-B Experiment Setup
Experiment Dataset
To ensure the accuracy of the experiment, we selected a balanced subset from the complete dataset to serve as the experimental dataset. The specific method is as follows: We chose 40 calligraphers and selected 40 characters that all of them have written. For each calligrapher, 90% of the characters were used for training, and the remaining characters were used for testing. This ensures that each test character is unseen during training for that particular calligrapher, while other characters by the same calligrapher and the same characters written by other calligraphers are used for training. This approach ensures that both the character shapes and the calligrapher’s style are adequately trained. Additionally, due to the varying number of characters for each calligrapher, we randomly selected up to four images for each character. Finally, we chose 12,985 images, including 11,689 images in tranining set and 1,296 images in test set. In this way, the negative impact of the long tail effect was successfully avoid .
Model Specifics
For the choice of VAE[16], we used the same pre-trained VAE as in LDM[17]. Specifically, the original image of $256\times 256\times 3$ size is mapped to a latent space of $32\times 32\times 4$ size. We set the number of iterations $N$ for our model to 4, with a hidden layer depth of 512, and the image segmentation patch size to 8. During training, we used a learning rate of $41e-$ 4 and trained on three A100 GPUs with a global batch size of 768. Ultimately, we selected the model at 288,000 steps (19,199 epochs) for subsequent experiment.
IV-C Evaluation Metrics
We evaluated our model from two perspectives: the structure and the style of the generated calligraphy. To assess the basic structural accuracy, we used Tencent’s handwritten OCR service [^2] https://cloud.tencent.com/document/product/866/36212 to recognize the generated calligraphy. Additionally, we used objective parameters such as IOU (Intersection over Union) and PSNR (Peak Signal-to-Noise Ratio) to measure the similarity between the generated calligraphy and the ground truth. A higher similarity indicates a closer match in style and a more reasonable structure. Lastly, we conducted a human evaluation experiment to verify how well our model captures the calligraphy style.
IV-D Experiment Results
Qualitative Evaluation
Tencent’s handwritten OCR service was used to recognize the generated calligraphy. The generated calligraphy were provided to the OCR API, which returned the recognized characters. We compared these with the character labels used to guide the generation process. If the recognized character matched the character label, we considered the generated character to have passed the test, indicating good structural integrity.
We randomly selected 60 images from each of the five different fonts for evaluation, considering that the recognition accuracy of handwritten OCR varies across different fonts. The results are shown in Table I. The models and datasets of other works have not been open-sourced, so this experiment was not conducted.
Font regular running cursive clerical seal
Moyun 0.783 0.583 0.167 0.117 0.03
TABLE I: OCR test result
The results indicate that our model performs well on commonly used fonts like regular script and running script. However, the recognition rates for less commonly used fonts such as cursive script, clerical script and seal script were lower. This is probably due to the significant differences between these fonts and modern simplified Chinese characters, which caused inaccuracies in OCR recognition. We will continue to evaluate structure of the generated calligraphy using other metrics.
We generated the same number of images as the entire test set using the prompts from the test set and evaluated the IOU (Intersection over Union). The results showed that our model significantly outperforms other models in terms of IOU. Using the same test set, we also evaluated the PSNR (Peak Signal-to-Noise Ratio) of the model. The results demonstrated that our model has a better performance in PSNR either. These findings are presented in Table II. This indicates that our model is better at fitting the structural integrity of calligraphy characters and more accurately replicating the style of calligraphy.
[…]
CalliGAN 0.325 -
ZiGAN 0.348 -
Zipeng Zhao[13] - 25.3734
Moyun(Ours) 0.810 32.0727
TABLE II: Qualitative measurements
Quantitative Evaluation
We designed a human evaluation experiment by selecting ten generated calligraphy samples and creating ten corresponding questions. Some of these calligraphy was shown in Figure 4. For each question, we provided one generated calligraphy, with four options: one calligraphy from the backbone, and three calligraphy from other calligraphers’ works. Both of them are the same character. The question posed was, “Please select the calligraphy that most closely matches the style of the given one.”
Figure 4: Generated calligraphy. Each column with different labels. The first row shows calligraphy generated by the model which were unseen before. and the second row is ground truth.
The questionnaire was distributed to 9 participants , and the results shows that $53.3$% generated calligraphy was paired with ground truth. This demonstrates that our model is capable of generating calligraphy in the specified style. The models and datasets of other works have not been open-sourced, so they were not included in the questionnaire.
V conclusion
In this paper, we proposed a new calligraphy generation model, “Moyun”, which could generate calligraphy in a specified style guided by the three labels: calligrapher, font, and character. The core idea was the introduction of Vision Mamba and the development of the TripleLabel control method. Additionally, we collected a large-scale, well-annotated, and properly binarized calligraphy dataset “Mobao”, which further demonstrated the effectiveness of our work.